The energy required to ionize magnesium is 738 kJ/mol . Part APart complete What minimum frequency of light is required to ionize magnesium?
Answer #1

Given:

Energy of 1 mol = 7.38*10^2 KJ/mol

= 7.38*10^5 J/mol

Find energy of 1 photon first

Energy of 1 photon = energy of 1 mol/Avogadro's number

= 7.38*10^5/(6.022*10^23)

= 1.226*10^-18 J

This is energy of 1 photon

Given:

E = 1.226*10^-18 J

use:

E = h*f

1.226*10^-18 J =(6.626*10^-34 J.s)*f

f = 1.85*10^15 Hz

Answer: 1.85*10^15 Hz